# From tug-of-war to Brownian Boost: explicit ODE solutions for player-funded stochastic-differential games

### v3

@@ 6.4. Equilibria converge to the putative Brownian Boost counterparts as $\kappa$ vanishes
$c(x)\,=\,2+\kappa\cdot 2\bigg(1-\frac{2\big(1+(1-\rho)x^{\rho}\big)}{(1+x^{\rho})^{2}}\bigg)\,+\,O(\kappa^{2})$ (61)
@@
$d(x)\,=\,2-\kappa\cdot 2\bigg(1-\frac{2}{(1+x^{\rho})^{2}}\Big((1-\rho)x^{\rho}+x^{2\rho}\Big)\bigg)\,+\,O(\kappa^{2})\,.$ (62)
@@
But $\frac{4\rho\,\beta^{1+\rho}}{(1+\beta^{\rho})^{2}}=\frac{4\rho\,x^{1+\rho}}{(1+x^{\rho})^{2}}+R$ where $|R|\leq 4\rho|\beta-x|D$, with $D=\sup\big\{\big|\tfrac{{\rm d}}{{\rm d}z}\tfrac{z^{1+\rho}}{(1+z^{\rho})^{2}}\big|:z\in(0,\infty)\big\}$. Hence,
$|\beta-x|\leq\frac{4\rho\,x^{1+\rho}}{(1+x^{\rho})^{2}}\,\kappa\;+\;4\rho|\beta-x|D\kappa+\beta\,|\Theta_{1}(\kappa,\rho,\beta)|\kappa^{2}\,$,
@@
$|\beta-x|\leq\bigg(\frac{4\rho\,x^{1+\rho}}{(1+x^{\rho})^{2}}\,\kappa\;+\;\beta\,|\Theta_{1}(\kappa,\rho,\beta)|\kappa^{2}\bigg)\big(1-4\rho D\kappa\big)^{-1}\,.$
Since $\beta\leq x$, we find that, provided that $\kappa\leq(8D)^{-1}$,
@@
Writing $v=u+1\geq 1$ and $D=(1-\kappa)v^{2}+2\kappa(1-\rho)v+2\kappa\rho$, we have $H^{\prime}(u)=4\kappa v\big((1-\rho)v+2\rho\big)D^{-2}$. Since $\kappa$ is at most a small positive constant, and $\rho\leq 1$, we see that $|H^{\prime}(u)|=O(\kappa)$ for $u\in[0,\infty)$.
@@
which since $\rho\leq 1$ is the desired asymptotic for $c$. The formula (62) for $d$ differs from that for $c$ in (61) only in a change $\rho\to-\rho$ in the linear-in-$\kappa$ coefficient in the denominator. The form of the estimates in the resulting proof are unaffected by this change, and the claimed $d$-asymptotic results. ∎
@@
Let $(\kappa,\rho)\in(0,1]^{2}$. Write $m_{i}(x)=m^{\rm def}_{i}(\kappa,\rho,x)$ (and use other like abbreviations) for the default solution. For $x\in D$ and $r\in\mathbb{R}$, we have that
$\kappa^{-1}m_{\lfloor\kappa^{-1}r\rfloor,\lfloor\kappa^{-1}r\rfloor+1}(x)$ $=$ $f_{\rho}(x,r)\big(1+\kappa E_{r}\big)\,\,\,\,\textrm{and}$ (63)
$\kappa^{-1}n_{\lfloor\kappa^{-1}r\rfloor+1,\lfloor\kappa^{-1}r\rfloor}(x)$ $=$ $g_{\rho}(x,r)\big(1+\kappa E_{r}\big)\,$,
where in each case the error $E_{r}$ is $O(1+x^{\rho})e^{25|r|}$. The quantities
$\kappa^{-1}M_{\lfloor\kappa^{-1}r\rfloor}=\kappa^{-1}m_{\lfloor\kappa^{-1}r\rfloor-1,\lfloor\kappa^{-1}r\rfloor+1}(x)$
@@
$\kappa^{-1}N_{\lfloor\kappa^{-1}r\rfloor}=\kappa^{-1}n_{\lfloor\kappa^{-1}r\rfloor-1,\lfloor\kappa^{-1}r\rfloor+1}(x)$
@@
$\kappa^{-2}a_{\lfloor\kappa^{-1}r\rfloor}(x)$ $=2\rho\;\frac{f_{\rho}(x,r^{\,1+\rho}\;g_{\rho}(x,r)^{\,\rho}}{\big(f_{\rho}(x,r)^{\rho}+g_{\rho}(x,r)^{\rho}\big)^{2}}\;\big(1+\kappa E_{r}\big),$ (64)
$\kappa^{-2}b_{\lfloor\kappa^{-1}r\rfloor}(x)$ $=2\rho\;\frac{f_{\rho}(x,r)^{\,\rho}\;g_{\rho}(x,r)^{\,1+\rho}}{\big(f_{\rho}(x,r)^{\rho}+g_{\rho}(x,r)^{\rho}\big)^{2}}\;\big(1+\kappa E_{r}\big),$ (65)
@@
Proof of Theorem 1.18(1). This is due to the estimates (64) and (65).
@@
$n_{k+1,k}(x)$ $=$ $\kappa\,x\cdot\prod_{i=0}^{k}\big(d_{i}(x)-1\big)\,,\,\,\,\textrm{and}$ (66)
@@
Adopt the shorthand $S(u)=S_{\rho}(x,u)$. It is straightforward that Proposition 6.3 implies that
$s_{\lfloor\kappa^{-1}u\rfloor}(x)\,=\,S(u)\,\Big(1+O\big(e^{(1+\rho)^{4}|u|}\big)\kappa\Big)\,.$ (67)
Apply the map $c$ to this relation, use Lemma 6.5(c) and that $\frac{x}{1+x^{2}}$ has a derivative that is uniformly bounded in absolute value to find that
$c\big(s_{\lfloor\kappa^{-1}u\rfloor}(x)\big)=c\big(S(u)\big)+O\big((1+S(u)^{\rho})e^{(1+\rho)^{4}|u|}\big)\kappa\,.$
The function $J(u)=S(u)^{\rho}$ solves the initial-value problem in Lemma 5.1. Since $|J^{\prime}|\leq 8\rho^{2}$ with $J(0)=x^{\rho}$, we have $J(u)\leq x^{\rho}e^{8\rho^{2}|u|}$ for $u\in\mathbb{R}$. Since $\rho\leq 1$, we obtain the naive upper bound on $S(u)^{\rho}$ of $x^{\rho}e^{8|u|}$; thus, the coefficient of $\kappa$ in the preceding display is $O(1+x^{\rho})e^{25|u|}$ for $u\in\mathbb{R}$.
@@
$c\big(s_{\lfloor\kappa^{-1}u\rfloor}(x)\big)-1\,=\,1+\kappa\cdot 2\bigg(1-\frac{2\big(1+(1-\rho)S(u)^{\rho}\big)}{(1+S(u)^{\rho})^{2}}\bigg)\,+\,E_{u}\,\kappa^{2}\,$, (68)
where $E_{u}=(1+x^{\rho})e^{25|u|}O(1)$. We see then that, for $r>0$,
$\prod_{i=0}^{\lfloor\kappa^{-1}r\rfloor}\big(c_{i}(x)-1\big)$ $=$ $\prod_{u\in\kappa\mathbb{Z}\cap[0,r]}\Bigg(1+\kappa\cdot 2\bigg(1-\frac{2\big(1+(1-\rho)S(u)^{\rho}\big)}{(1+S(u)^{\rho})^{2}}\bigg)+E_{u}\,\kappa^{2}\Bigg)$
$=$ $\exp\Bigg\{2\kappa\sum_{u\in\kappa\mathbb{Z}\cap[0,r]}\bigg(1-\frac{2\big(1+(1-\rho)S(u)^{\rho}\big)}{(1+S(u)^{\rho})^{2}}\bigg)\,+\,\kappa\,E_{r}\Bigg\}$
$=$ $\exp\Bigg\{2\int_{0}^{r}\bigg(1-\frac{2\big(1+(1-\rho)S(u)^{\rho}\big)}{(1+S(u)^{\rho})^{2}}\bigg)\,{\rm d}u\Bigg\}\Big(1+\kappa E_{r}\Big)\,$,
where the error terms $E$ may differ from line to line, subject to the condition given when they were introduced above. Since the exponential expression in the final line equals $f_{\rho}(x,r)$, we obtain the sought bound on $\kappa^{-1}m_{\lfloor\kappa^{-1}r\rfloor,\lfloor\kappa^{-1}r\rfloor+1}(x)$ for $r>0$. And also when $r<0$, provided that the product and sum expressions in the preceding display are interpreted compatibly with the convention for negatively indexed products in (1.15).
@@
$d\big(s_{\lfloor\kappa^{-1}u\rfloor}(x)\big)-1\,=\,1-\kappa\cdot 2\bigg(1-\frac{2\big(1-\rho)S(u)^{\rho}+S(u)^{2\rho}\big)}{(1+S(u)^{\rho})^{2}}\bigg)\,+\,E_{u}\,\kappa^{2}\,$,
@@
$\prod_{i=0}^{\lfloor\kappa^{-1}r\rfloor}\big(d_{i}(x)-1\big)\,=\,\exp\Bigg\{-2\int_{0}^{r}\bigg(1-\frac{2\big((1-\rho)S(u)^{\rho}+S(u)^{2\rho}\big)}{(1+S(u)^{\rho})^{2}}\bigg)\,{\rm d}u\Bigg\}\Big(1+\kappa E_{r}\Big)$ (69)
Noting the factor of $x$ on the right-hand side of (66), we multiply (69) by $x$ and note that the resulting right-hand term $x\cdot\exp\{-2I\}$ equals $g_{\rho}(x,r)$. The bound on $\kappa^{-1}n_{\lfloor\kappa^{-1}r\rfloor+1,\lfloor\kappa^{-1}r\rfloor}(x)$ follows.
To obtain the assertion made in regard to $\kappa^{-1}M_{\lfloor\kappa^{-1}r\rfloor}$, sum (63) for values $r-\kappa$ and $r$, and use the differentiability of $f_{\rho}(x,r)$ at $r$ to absorb via a factor of $1+O(\kappa)$ the error arising from the microscopic unit index displacement. (The derivative in $r$ is readily seen to be bounded on compact subsets of $\mathbb{R}$; in fact, decay at infinity means that this is true on all of $\mathbb{R}$. So the implied constant in the $O(\kappa)$ term may be chosen independently of $r\in\mathbb{R}$.)
Likewise for $\kappa^{-1}N_{\lfloor\kappa^{-1}r\rfloor}$. Applying these estimates to the formulas for $a_{i}$ and $b_{i}$ in terms of $M_{i}$ and $N_{i}$ in Proposition 6.1(1), we obtain the stated asymptotics for $a_{i}$ and $b_{i}$ and thus complete the proof of Proposition 6.6. ∎
@@ 6.5. The low $\kappa$ limit of $\lambda_{\rm max}$
Proof of Theorem 1.22. In light of reduction to standard solutions by basic symmetries, Definition 2.8 and Proposition 2.9(4), it suffices to show that there exist positive $C$ and $c$ such that, for $\kappa$ small enough,
$\left|\,\frac{n_{\infty,-\infty}}{m_{-\infty,\infty}}-1\,\right|\,\leq\,C\kappa^{c}$ (70)
@@
Write $x=\phi_{0}$. The plan is to argue that $m_{-\infty,\infty}$ equals $m_{-1,0}\int_{\mathbb{R}}f_{\rho}(x,u)\,{\rm d}u$ up to an error that vanishes as $\kappa\searrow 0$, and that $n_{\infty,-\infty}$ similarly approximates $m_{-1,0}\int_{\mathbb{R}}g_{\rho}(x,u)\,{\rm d}u$. The integrals are equal by Proposition 5.4(5), as desired. To implement this approach, we will use the approximations of $\kappa$-scaled $m$- and $n$-differences by $f_{\rho}(x,u)$ and $g_{\rho}(x,u)$ found in Proposition 6.6. These approximations worsen for indices that are high multiples of $\kappa^{-1}$ because the mimicry of the $s$-orbit by the $S_{\rho}$-flow (as gauged by Proposition 6.3) may have deteriorated. So we will attempt the comparison only on a short scale, delimited by a continuous-time parameter $z$. We handle the longer scale via the next result.
@@
[…]
@@
where $z=c\log\kappa^{-1}$.
Proof. Maintain the shorthand $S(u)=S_{\rho}(x,u)$. It is also useful to have compact notation for the scaled $s$-iterates, and we set $S_{\kappa}(u)=s_{\lfloor\kappa^{-1}u\rfloor}(x)$ with $x=\phi_{0}$. Since $z=c\log\kappa^{-1}$, the bound $S_{\kappa}(u)\leq 2S(z)$ holds for $u=z$ by (67), provided that we make a suitably small choice of $c>0$. And since $s$ is sub-diagonal by Lemma 2.3(3), this bound also holds for all $u\geq z$.
@@
$c\big(S_{\kappa}(u)\big)-1=1+2\,\Phi_{f}\big(S_{\kappa}(u)\big)\kappa+E_{u}\kappa^{2}$
@@
Note that $S$ solves the differential equation in Definition 1.2, so $S(y)\to 0$ as $y\to\infty$. Note also that $\lim_{y\searrow 0}\Phi_{f}(y)=-1$. For a suitably small choice of $\kappa_{0}=\kappa_{0}(c)$, the condition $\kappa\in(0,\kappa_{0})$ thus ensures that $S(z)<\epsilon$ where $\epsilon>0$ is such that $\Phi_{f}(y)\leq-3/4$ for $y\in(0,2\epsilon)$. Since $S_{\kappa}(u)\leq 2S(z)$ for $u\geq z$, the linear coefficient in the last display is at most 3/2-3/2. Since $E_{u}\kappa^{2}\leq\kappa/2$ by choosing $\kappa_{0}$ suitably, we see that $c(S_{\kappa}(u))-1\leq 1-\kappa$ for $u\geq z$.
By taking a ratio of equalities of the form (13), we obtain
$m_{\lfloor z\kappa^{-1}\rfloor+i,\lfloor z\kappa^{-1}\rfloor+i+1}m_{\lfloor z\kappa^{-1}\rfloor,\lfloor z\kappa^{-1}\rfloor+1}^{-1}=\prod_{j=0}^{i-1}\Big(c\big(s_{j}\big(S_{\kappa}(z)\big)\big)-1\Big)$
since this ratio of $m$-differences coincides with this ratio for the default solution with the same value of $\phi_{0}$. Noting also that $s_{j}\big(S_{\kappa}(z)\big)=S_{\kappa}(z+\kappa j)$, we find from $m_{\lfloor z\kappa^{-1}\rfloor,\infty}=\sum_{j=\lfloor z\kappa^{-1}\rfloor}^{\infty}m_{j,j+1}$ that
$m_{\lfloor z\kappa^{-1}\rfloor,\infty}\,=\,m_{\lfloor z\kappa^{-1}\rfloor,\lfloor z\kappa^{-1}\rfloor+1}\sum_{i=0}^{\infty}\prod_{j=0}^{i-1}\big(c(S_{\kappa}(z+\kappa j))-1\big)\,\leq\,m_{\lfloor z\kappa^{-1}\rfloor,\lfloor z\kappa^{-1}\rfloor+1}\cdot\kappa^{-1}\,$, (71)
where in the final bound, we applied the just obtained upper bound of $1-\kappa$ on $c(S_{\kappa}(u))-1$.

### v4

@@ 6.4. Equilibria converge to the putative Brownian Boost counterparts as $\kappa$ vanishes
$c(x)\,=\,2+\kappa\cdot 2\bigg(1-\frac{2\big(1+(1-\rho)x^{\rho}\big)}{(1+x^{\rho})^{2}}\bigg)\,+\,O(\kappa^{2})$ (6.10)
@@
$d(x)\,=\,2-\kappa\cdot 2\bigg(1-\frac{2}{(1+x^{\rho})^{2}}\Big((1-\rho)x^{\rho}+x^{2\rho}\Big)\bigg)\,+\,O(\kappa^{2})\,.$ (6.11)
@@
But $\frac{4\rho\,\beta^{1+\rho}}{(1+\beta^{\rho})^{2}}=\frac{4\rho\,x^{1+\rho}}{(1+x^{\rho})^{2}}+R$ where $|R|\leq 4\rho|\beta-x|L$, with $L=\sup\big\{\big|\tfrac{{\rm d}}{{\rm d}z}\tfrac{z^{1+\rho}}{(1+z^{\rho})^{2}}\big|:z\in(0,\infty)\big\}$. Hence,
$|\beta-x|\leq\frac{4\rho\,x^{1+\rho}}{(1+x^{\rho})^{2}}\,\kappa\;+\;4\rho|\beta-x|L\kappa+\beta\,|\Theta_{1}(\kappa,\rho,\beta)|\kappa^{2}\,$,
@@
$|\beta-x|\leq\bigg(\frac{4\rho\,x^{1+\rho}}{(1+x^{\rho})^{2}}\,\kappa\;+\;\beta\,|\Theta_{1}(\kappa,\rho,\beta)|\kappa^{2}\bigg)\big(1-4\rho L\kappa\big)^{-1}\,.$
Since $\beta\leq x$, we find that, provided that $\kappa\leq(8L)^{-1}$,
@@
Writing $v=u+1\geq 1$ and $\Delta=(1-\kappa)v^{2}+2\kappa(1-\rho)v+2\kappa\rho$, we have $H^{\prime}(u)=4\kappa v\big((1-\rho)v+2\rho\big)\Delta^{-2}$. Since $\kappa$ is at most a small positive constant, and $\rho\leq 1$, we see that $|H^{\prime}(u)|=O(\kappa)$ for $u\in[0,\infty)$.
@@
which since $\rho\leq 1$ is the desired asymptotic for $c$. The formula (6.11) for $d$ differs from that for $c$ in (6.10) only in a change $\rho\to-\rho$ in the linear-in-$\kappa$ coefficient in the denominator. The form of the estimates in the resulting proof are unaffected by this change, and the claimed $d$-asymptotic results. ∎
@@
Let $(\kappa,\rho)\in(0,1]^{2}$. Write $m_{i}(x)=m^{\rm def}_{i}(\kappa,\rho,x)$ (and use other like abbreviations) for the default solution. For $\xi\in D$ and $u\in\mathbb{R}$, we have that
$\kappa^{-1}m_{\lfloor\kappa^{-1}u\rfloor,\lfloor\kappa^{-1}u\rfloor+1}(\xi)$ $=$ $f_{\rho}(\xi,u)\big(1+\kappa E_{u}\big)\,\,\,\,\textrm{and}$ (6.12)
$\kappa^{-1}n_{\lfloor\kappa^{-1}u\rfloor+1,\lfloor\kappa^{-1}u\rfloor}(\xi)$ $=$ $g_{\rho}(\xi,u)\big(1+\kappa E_{u}\big)\,$,
where in each case the error $E_{u}$ is $O(1+\xi^{\rho})e^{C_{1}|u|}$; here, the constant $C_{1}$ is inherited from Proposition 6.3 via the relation $\stackrel{{\scriptstyle\kappa}}{{\simeq}}$. The quantities
$\kappa^{-1}M_{\lfloor\kappa^{-1}u\rfloor}=\kappa^{-1}m_{\lfloor\kappa^{-1}u\rfloor-1,\lfloor\kappa^{-1}u\rfloor+1}(\xi)$
@@
$\kappa^{-1}N_{\lfloor\kappa^{-1}u\rfloor}=\kappa^{-1}n_{\lfloor\kappa^{-1}u\rfloor-1,\lfloor\kappa^{-1}u\rfloor+1}(\xi)$
@@
$\kappa^{-2}a_{\lfloor\kappa^{-1}u\rfloor}(\xi)$ $=2\rho\;\frac{f_{\rho}(\xi,u)^{\,1+\rho}\;g_{\rho}(\xi,u)^{\,\rho}}{\big(f_{\rho}(\xi,u)^{\rho}+g_{\rho}(\xi,u)^{\rho}\big)^{2}}\;\big(1+\kappa E_{u}\big),$ (6.13)
$\kappa^{-2}b_{\lfloor\kappa^{-1}u\rfloor}(\xi)$ $=2\rho\;\frac{f_{\rho}(\xi,u)^{\,\rho}\;g_{\rho}(\xi,u)^{\,1+\rho}}{\big(f_{\rho}(\xi,u)^{\rho}+g_{\rho}(\xi,u)^{\rho}\big)^{2}}\;\big(1+\kappa E_{u}\big),$ (6.14)
@@
Proof of Theorem 1.18(1). This is due to the estimates (6.13) and (6.14).
@@
$n_{k+1,k}(x)$ $=$ $\kappa\,x\cdot\prod_{i=0}^{k}\big(d_{i}(x)-1\big)\,,\,\,\,\textrm{and}$ (6.15)
@@
We now set $x$ equal to the element $\xi\in D$ given in the proposition; we also adopt the shorthand $S(u)=S_{\rho}(\xi,u)$. It is straightforward that Proposition 6.3 implies that
$s_{\lfloor\kappa^{-1}u\rfloor}(\xi)\,=\,S(u)\,\Big(1+O\big(e^{C_{1}|u|}\big)\kappa\Big)\,;$ (6.16)
the exponential growth has a uniform rate since $\xi$ lies in the central domain $D$, where it is bounded away from zero and infinity. Apply the map $c$ to this relation, use [^1] Recall the convention for result references stated after (3.2) Lemma 6.5(c) and that $\frac{z^{\rho}}{(1+z^{\rho})^{2}}$ has a derivative that is uniformly bounded in absolute value to find that
$c\big(s_{\lfloor\kappa^{-1}u\rfloor}(\xi)\big)=c\big(S(u)\big)+O\big((1+S(u)^{\rho})e^{C_{1}|u|}\big)\kappa\,.$
The function $J(u)=S(u)^{\rho}$ solves the initial-value problem in Lemma 5.1. Since $|J^{\prime}|\leq 8\rho^{2}$ with $J(0)=\xi^{\rho}$, we have $J(u)\leq\xi^{\rho}e^{8\rho^{2}|u|}$ for $u\in\mathbb{R}$. Since $\rho\leq 1$, we obtain the naive upper bound on $S(u)^{\rho}$ of $\xi^{\rho}e^{8|u|}$; absorbing the additive eight by an increase in the value of $C_{1}$, the coefficient of $\kappa$ in the preceding display may thus be written $O(1+\xi^{\rho})e^{C_{1}|u|}$ for $u\in\mathbb{R}$.
@@
$c\big(s_{\lfloor\kappa^{-1}u\rfloor}(\xi)\big)-1\,=\,1+\kappa\cdot 2\bigg(1-\frac{2\big(1+(1-\rho)S(u)^{\rho}\big)}{(1+S(u)^{\rho})^{2}}\bigg)\,+\,E_{u}\,\kappa^{2}\,$, (6.17)
where $E_{u}=(1+\xi^{\rho})e^{C_{1}|u|}O(1)$. We see then that, for $u>0$,
$\prod_{i=0}^{\lfloor\kappa^{-1}u\rfloor}\big(c_{i}(\xi)-1\big)$ $=$ $\prod_{r\in\kappa\mathbb{Z}\cap[0,u]}\Bigg(1+\kappa\cdot 2\bigg(1-\frac{2\big(1+(1-\rho)S(r)^{\rho}\big)}{(1+S(r)^{\rho})^{2}}\bigg)+E_{r}\,\kappa^{2}\Bigg)$
$=$ $\exp\Bigg\{2\kappa\sum_{r\in\kappa\mathbb{Z}\cap[0,u]}\bigg(1-\frac{2\big(1+(1-\rho)S(r)^{\rho}\big)}{(1+S(r)^{\rho})^{2}}\bigg)\,+\,\kappa\,E_{u}\Bigg\}$
$=$ $\exp\Bigg\{2\int_{0}^{u}\bigg(1-\frac{2\big(1+(1-\rho)S(r)^{\rho}\big)}{(1+S(r)^{\rho})^{2}}\bigg)\,{\rm d}r\Bigg\}\Big(1+\kappa E_{u}\Big)\,$,
the second equality due to the standard estimate $\prod(1+t_{i})=\exp\big\{\sum t_{i}+O(\sum t_{i}^{2})\big\}$. The error terms $E$ may differ from line to line, subject to the condition given when they were introduced above. Since the exponential expression in the final line equals $f_{\rho}(\xi,u)$, we obtain the sought bound on $\kappa^{-1}m_{\lfloor\kappa^{-1}u\rfloor,\lfloor\kappa^{-1}u\rfloor+1}(\xi)$ for $u>0$. And also when $u<0$, provided that the product and sum expressions in the preceding display are interpreted compatibly with the convention for negatively indexed products in (1.10).
@@
$d\big(s_{\lfloor\kappa^{-1}u\rfloor}(\xi)\big)-1\,=\,1-\kappa\cdot 2\bigg(1-\frac{2\big((1-\rho)S(u)^{\rho}+S(u)^{2\rho}\big)}{(1+S(u)^{\rho})^{2}}\bigg)\,+\,E_{u}\,\kappa^{2}\,$,
@@
$\prod_{i=0}^{\lfloor\kappa^{-1}u\rfloor}\big(d_{i}(\xi)-1\big)\,=\,\exp\Bigg\{-2\int_{0}^{u}\bigg(1-\frac{2\big((1-\rho)S(r)^{\rho}+S(r)^{2\rho}\big)}{(1+S(r)^{\rho})^{2}}\bigg)\,{\rm d}r\Bigg\}\Big(1+\kappa E_{u}\Big)\,.$ (6.18)
Noting the factor of $\xi$ on the right-hand side of (6.15), we multiply (6.18) by $\xi$ and find that the resulting right-hand term $\xi\cdot\exp\{-2I\}$ equals $g_{\rho}(\xi,u)$. The bound on $\kappa^{-1}n_{\lfloor\kappa^{-1}u\rfloor+1,\lfloor\kappa^{-1}u\rfloor}(\xi)$ follows.
To obtain the assertion made in regard to $\kappa^{-1}M_{\lfloor\kappa^{-1}u\rfloor}$, sum (6.12) for values $u-\kappa$ and $u$, and use the differentiability of $f_{\rho}(\xi,u)$ at $u$ to absorb via a factor of $1+O(\kappa)$ the error arising from the microscopic unit index displacement. (The derivative in $u$ is readily seen to be bounded on compact subsets of $\mathbb{R}$; in fact, decay at infinity means that this is true on all of $\mathbb{R}$. So the implied constant in the $O(\kappa)$ term may be chosen independently of $u\in\mathbb{R}$.)
Likewise for $\kappa^{-1}N_{\lfloor\kappa^{-1}u\rfloor}$. Applying these estimates to the formulas for $a_{i}$ and $b_{i}$ in terms of $M_{i}$ and $N_{i}$ in Proposition 6.1(1), we obtain the stated asymptotics for $a_{i}$ and $b_{i}$ and thus complete the proof of Proposition 6.6. ∎
@@ 6.5. The low $\kappa$ limit of $\lambda_{\rm max}$
Proof of Theorem 1.23. In light of reduction to standard solutions by basic symmetries, Definition 2.8 and Proposition 2.9(4), it suffices to show that there exist positive $C$ and $\eta$ such that, for $\kappa$ small enough,
$\left|\,\frac{n_{\infty,-\infty}}{m_{-\infty,\infty}}-1\,\right|\,\leq\,C\kappa^{\eta}$ (6.19)
@@
Write $x=\phi_{0}$. The plan is to argue that $m_{-\infty,\infty}$ equals $m_{-1,0}\int_{\mathbb{R}}f_{\rho}(x,r)\,{\rm d}r$ up to an error that vanishes as $\kappa\searrow 0$, and that $n_{\infty,-\infty}$ similarly approximates $m_{-1,0}\int_{\mathbb{R}}g_{\rho}(x,r)\,{\rm d}r$. The integrals are equal by Proposition 5.4(5), as desired. To implement this approach, we will use the approximations of $\kappa$-scaled $m$- and $n$-differences by $f_{\rho}(x,u)$ and $g_{\rho}(x,u)$ found in Proposition 6.6. These approximations worsen for indices that are high multiples of $\kappa^{-1}$ because the mimicry of the $s$-orbit by the $S_{\rho}$-flow (as gauged by Proposition 6.3) may have deteriorated. So we will attempt the comparison only on a short scale, delimited by a continuous-time parameter $z$ of the form $\eta\log\kappa^{-1}$. We handle the longer scale via the next result.
@@
There exist positive $\eta$, $c_{0}$ and $C$ such that, for $(a,b,m,n)\in$ ABMN($\kappa,\rho$) with $\phi_{0}\in D$,
@@
where $z=\eta\log\kappa^{-1}$.
In the argument for Theorem 1.23, the flow $S_{\rho}$ acts as an analytic proxy for the discrete dynamics. So we set the index $\xi$ that specifies the flow $S_{\rho}(\xi,\bullet):\mathbb{R}\to(0,\infty)$ equal to the given initial ratio $x=\phi_{0}$, and write $S(u)=S_{\rho}(x,u)$ for the resulting flow. We also denote the scaled $s$-iterates that the flow tracks by $\mathsf{S}_{\kappa}(u):=s_{\lfloor\kappa^{-1}u\rfloor}(x)$.
Proof of Lemma 6.7. Since $z=\eta\log\kappa^{-1}$, the bound $\mathsf{S}_{\kappa}(u)\leq 2S(z)$ holds for $u=z$ by (6.16), provided that we make a suitably small choice of $\eta>0$. And since $s$ is sub-diagonal by Lemma 2.1(3), this bound also holds for all $u\geq z$.
@@
$c\big(\mathsf{S}_{\kappa}(u)\big)-1=1+2\,\Phi_{f}\big(\mathsf{S}_{\kappa}(u)\big)\kappa+E_{u}\kappa^{2}$
@@
Note that $S$ solves the differential equation in Definition 1.14, so $S(y)\to 0$ as $y\to\infty$. Note also that $\lim_{y\searrow 0}\Phi_{f}(y)=-1$. For a suitably small choice of $\kappa_{0}=\kappa_{0}(\eta)$, the condition $\kappa\in(0,\kappa_{0})$ thus ensures that $S(z)<\epsilon$ where $\epsilon>0$ is such that $\Phi_{f}(y)\leq-3/4$ for $y\in(0,2\epsilon)$. Since $\mathsf{S}_{\kappa}(u)\leq 2S(z)$ for $u\geq z$, the linear coefficient in the last display is at most 3/2-3/2. Since $E_{u}\kappa^{2}\leq\kappa/2$ by choosing $\kappa_{0}$ suitably, we see that $c(\mathsf{S}_{\kappa}(u))-1\leq 1-\kappa$ for $u\geq z$.
By taking a ratio of equalities of the form (1.12), we obtain
$m_{\lfloor z\kappa^{-1}\rfloor+i,\lfloor z\kappa^{-1}\rfloor+i+1}m_{\lfloor z\kappa^{-1}\rfloor,\lfloor z\kappa^{-1}\rfloor+1}^{-1}=\prod_{j=0}^{i-1}\Big(c\big(s_{j}\big(\mathsf{S}_{\kappa}(z)\big)\big)-1\Big)$
since this ratio of $m$-differences coincides with this ratio for the default solution with the same value of $\phi_{0}$. Certainly the right-hand side is a product of positive quantities: see the discussion leading up to (3.2) in Section 3.1. Noting also that $s_{j}\big(\mathsf{S}_{\kappa}(z)\big)=\mathsf{S}_{\kappa}(z+\kappa j)$, we find from $m_{\lfloor z\kappa^{-1}\rfloor,\infty}=\sum_{j=\lfloor z\kappa^{-1}\rfloor}^{\infty}m_{j,j+1}$ that
$m_{\lfloor z\kappa^{-1}\rfloor,\infty}\,=\,m_{\lfloor z\kappa^{-1}\rfloor,\lfloor z\kappa^{-1}\rfloor+1}\sum_{i=0}^{\infty}\prod_{j=0}^{i-1}\big(c(\mathsf{S}_{\kappa}(z+\kappa j))-1\big)\,\leq\,m_{\lfloor z\kappa^{-1}\rfloor,\lfloor z\kappa^{-1}\rfloor+1}\cdot\kappa^{-1}\,$, (6.20)
where in the final bound, we applied the just obtained upper bound of $1-\kappa$ on $c(\mathsf{S}_{\kappa}(u))-1$.